# Uncertainty quantification for intervals

### v1

@@ 4 Simulation studies
$F(t,x)=P\left(T\leq t\mid X=x\right)=P\left(p\log\left(sT\right)\leq p\log\left(st\right)\mid X=x\right)$
$=P\left(-r(x)+H\leq p\log\left(st\right)\right)$
$=P\left(H\leq p\log\left(st\right)+r(x)\right)$
@@
Therefore, Equation 6 hides a Cox model with baseline cumulative hazard $H_{0}(t)=(st)^{p}$, corresponding to a Weibull distribution of scale and shape parameters $s$ and $p$ respectively. We could have changed $p\log\left(sT\right)$ by any strictly monotonically increasing function (first, we just need to preserve inequalities inside probabilities and second, inverse of these functions is always guaranteed to exist).
4.1 Verfication of asympotic properties
We empirically corroborate what Theorem 3.6 implies: the probability that the new observation $T_{N+1}$ falls within the prediction interval $\widehat{\mathcal{C}}_{(1-\alpha)}^{*}(X_{N+1})$ is approximately $1-\alpha$ as the sample size grows. The term $o_{\mathbb{P}}(1)$ denotes a small error term that converges to zero in probability as the sample size increases, signifying that the approximation becomes more accurate with larger sample sizes.
In this setup, uncervals is working under $e=1$ operating mode, i.e. we uniformly randomize $\Phi^{*}$ over the interval whose upper and lower borders are $\hat{F}_{1}$ evaluations on the original intervals.
We prepare four different simulation scenarios, the setup of which is described in Table 1. The underlying true times are simulated according to model 6. These are then censored with a case II interval-censoring mechanism (see https://cran.r-project.org/web/packages/icenReg/index.html). Time between inspections is distributed as runif(min = 0, max = inspectLength). Then, we check the frequency with which the intervals outputted by our algorithm contain the true time-to-event over 100 test points. We average the results across 100 different training sets. This procedure is repeated for six different nominal coverages ($0.5,0.7,0.8,0.9,0.95,0.99$) and four ($100,200,500,1000$) different sample sizes. For each row in Table 1, we choose different base estimators for $\hat{F}_{1}$. In the “Linear PH” case, we use ic_sp(Surv(L, U, type = ’interval2’), model = ’ph’) from the icenReg package to fit $\hat{F}_{1}$. In the “Non-linear PH” case, we use Interval Censored Recursive Forests [33] with ntree=20, nfold=2. The marginal case is the same as “Linear PH” but with no covariates specified. For the “Oracle” mode, the true survival function, as derived in LABEL:ors, is directly passed to our approach. The results are visible in Figure 4.
@@ 4.1 Verfication of asympotic properties
((a))
((b))
((c))
((d))
Figure 4: Empirical vs nominal coverage for the setup in Section 4.1. It must be noted that throughout our simulations, uncervals was operating $b=0.5$. This means that we are considering intervals whose image by $S(\cdot,x)$ is centered around 0.5.
4.2 Comparison of different LPBs
We consider three different sample sizes and numbers of covariates. For each sample size and number of covariates, we evaluate three different algorithms for uncertainty quantification. We propose $\alpha_{1}=0.8$, $\alpha_{2}=0.9$, $\alpha_{3}=0.95$. In all cases, we set $b=1$ (therefore LPB).
  1. 1.
    Checking if $\hat{F}_{1}(t,X_{\text{new}})$ is greater than $\alpha$: the LPB has as a lower bound the estimated $\alpha$ quantile (naive quantile).
  2. 2.
    Our approach with $e=0$, this is, considering evaluations of $\hat{F}_{1}$ just on the left borders, thus the trivial extension of [5]. Instead of randomizing $\Phi_{i}^{*}\leftarrow\texttt{runif}\left(\hat{F}_{1}\left(L_{j},X_{j}\right),\hat{F}_{1}\left(U_{j},X_{j}\right)\right)$ we set $\Phi_{i}^{*}\leftarrow\hat{F}_{1}\left(L_{j},X_{j}\right)$ so that the finite sample coverage bound holds.
  3. 3.
    Our approach with $e=*$ (randomizing between border evaluations).
The results are visible in Tables 2-10.
Table 2: Results for naive quantile and Level $1-\alpha_{1}$
@@ 4.2 Comparison of different LPBs
Table 3: Results for uncervals, e=0 and Level $1-\alpha_{1}$
@@
Table 4: Results for uncervals, e=* and Level $1-\alpha_{1}$
@@
Table 5: Results for naive quantile and Level $1-\alpha_{2}$
@@
Table 6: Results for uncervals, e=0 and Level $1-\alpha_{2}$
@@
Table 7: Results for uncervals, e=* and Level $1-\alpha_{2}$
@@
Table 8: Results for naive quantile and Level $1-\alpha_{3}$
@@
Table 9: Results for uncervals, e=0 and Level $1-\alpha_{3}$
@@
Table 10: Results for uncervals, e=* and Level $1-\alpha_{3}$
@@
4.3 Evaluating Conformalized Survival Analysis
In the Conformalized Survival Analysis paper [5], the censoring mechanism is leveraged to mitigate the conservativeness of naive approaches driven by small censoring times. By introducing a threshold $c_{0}$ and focusing on subpopulations where $C\geq c_{0}$, they address the distributional shift between subpopulations and the whole population. This shift arises because patients with larger censoring times tend to be healthier, leading to different joint and conditional distributions of the variables $X$, $C$, and $T$. The authors suggest using a secondary censored outcome $T\wedge c_{0}$ and highlight that while there is a covariate shift, this can be adjusted by reweighting the samples through weighted conformal inference [65], allowing for a calibrated lower prediction bound (LPB) on $T\wedge c_{0}$ and thus on $T$. This approach reduces the power loss due to censoring, but depends on choosing the threshold hyperparameter $c_{0}$. We performed a sensitivity analysis of the cfsurv function from the cfsurvival package. Performance in terms of empirical coverage depends strongly on the chosen $c_{0}$, as seen in Figure 2. We simulate as in Section 4 but now setting par_shape=2, par_scale=1, par_inspections=5, $n=2000$, $X\sim U(0,2)$ and $r(X)=X$ without left censoring. Under this configuration, right-censoring amounts to approximately $30\%$. For a grid of $c_{0}$ ranging from $0.1$ to $1$ we launch conformalized survival analysis by setting $\tilde{T}$ to the lower bounds, the event indicator to $1\{U=+\infty\}$ and the censoring times to $1\{U=+\infty\}L+1\{U<+\infty\}U$. See also in Figure 2 how empirical coverage depends on the hyperparameter $c_{0}$. $\alpha$ was set to $0.1$. Also, recall that the methodology in [5] requires observing $C$ even when survival time $T$ is available.
5 Evaluation and comparision of the models with empirical data
5.1 Systolic Blood Pressure Intervention Trial (SPRINT)
In order to assess the conservative non-asymptotic generalization for interval-censored data of the right-censored algorithm provided in [5], in comparison with our algorithm, we the sleep time from the NIH’s Systolic Blood Pressure Intervention Trial (SPRINT) study [27] as an outcome, in which we manage to recover the censoring time $C>0$ as it is needed in [5].
SPRINT [27] is a clinical trial that was conducted with the goal to inform the new blood pressure medication guidelines in the US population by testing the effects that a lower blood pressure target has on reducing heart disease risk. Observational studies had shown that individuals with lower systolic blood pressure (SBP) levels had fewer complications and deaths due to cardiovascular disease (CVD). Building on this observation, SPRINT was designed to test the effects of a lower blood pressure target on reducing heart disease risk. Specifically, SPRINT aimed to compare treating high blood pressure to a target SBP goal of less than $120$ mmHg against treating to a goal of less than $140$ mmHg.
The “Ambulatory Blood Pressure Monitoring” (ABPM) study, an ancillary study within the SPRINT trial, investigates the effects of different hypertension management strategies on nighttime blood pressure. This research explores whether targeting a clinic systolic blood pressure (SBP) of less than $120$ mmHg provided clinical advantages in reducing nighttime BP, a critical predictor of cardiovascular events, over a standard target of less than $140$ mmHg.
In our analysis, we propose examining the relationship between “Time that sleep ended” and “Morning Systolic BP ”. This relationship is pertinent because morning BP can influence wake-up times due to the physiological stress or discomfort associated with elevated BP levels. Understanding how morning systolic BP affects when individuals wake up can provide valuable insights into the interplay between BP management and sleep quality. By analyzing this connection, we aim to enhance hypertension treatment strategies, ensuring better BP control throughout the night and improving overall health outcomes for patients.
We build a couple of semi-synthetic responses: an interval-censored one employable by our algorithm and a right-censored one that can be fed to [5]. Assume that people in the study misunderstood their indications and did not write down the time when they woke up. However, we have access to the time that was set in their alarms and they remember if they woke up naturally or using the alarm. We craft the interval-censored $(L,U)$ and the right-censored $(\widetilde{T},\textrm{Event indicator})$ following Table 11.
L U $\mathbf{\tilde{T}}$ Event indicator C T
Natural wake up (zzzzz)--! ( ! ( 1 ! )
Alarm wake up (zz!--) ! +$\infty$+ ∞ ! 0 ! )
Table 11: Casuistic that has been followed in order to build the semi-synthetic data. The original interval-censored observations were given by two numbers ( and ) representing the time when sleep started and ended respectively. ! encodes an artificial independent censoring time that has been simulated according to rexp(ninitial, rate = 0.01), ensuring that the new target amounts to 10% right-censoring approximately.
The final dataset analyzed contains a total of $863$ individuals, which we randomly split into 80% train and 20% test in two independent datasets.
On the one hand, we consider the approach in [5] using a Cox model to fit the conditional survival function and the corresponding quantile conformal score and we feed it with its semi-synthetic training data, which are samples from $(X,\widetilde{T},\textrm{Event indicator},C)$ where $X$ is MORNING_SYSTOLIC. Since $\tilde{T}\leq T$, any calibrated lower predictive bound on the censored survival time $\tilde{T}$ is also a calibrated LPB on the uncensored survival time $T$. We try several values of $c_{0}$ and the coverage is always 1, a consequence of overconservative behavior.
On the other hand, we challenge our approach using as a base model ic_sp(Surv(L, U, type = ’interval2’), model = ’ph’) to fit $\hat{F}_{1}$. We repeat the procedure $B=100$ times by randomizing the train-test split and, within the training set, the split used in the two phases of conformal inference. We obtain a median coverage of $0.8576$, indicating that the Cox model might be misspecified in this setting. Next, we try Interval Censored Recursive Forests another $B=100$ times and observe a median coverage equal to $0.93$.
5.2 NHANES Physical Activity Example
The National Health and Nutrition Examination Survey (NHANES) is a comprehensive program conducted by the National Center for Health Statistics (NCHS), with the primary objective of collecting health and nutrition data from the United States population. In this study, we utilized a subset of $n=2977$ patients aged between 50 and 80 years, gathered from NHANES waves 2003-2006. This subset contains high-resolution physical activity information measured by means of accelerometer devices during 3-7 days with a time resolution of one physical activity measure per minute.
From a public health and epidemiological perspective, establishing a statistical association between physical activity and survival in the elderly population is of paramount importance. This analysis can serve as the foundation for promoting healthy habits across the population and designing targeted interventions to mitigate the adverse effects of inactivity and the natural decline in functional capacity that occurs with aging [43, 81]. Furthermore, given that physical inactivity is a modifiable risk factor for various chronic diseases, including heart disease, diabetes, and specific types of cancer, the implementation of effective physical activity policies has the potential to extend the lifespan of this at-risk population, enhance their quality of life, and significantly reduce healthcare costs while alleviating the burden on the medical healthcare systems.
There is a corpus of prior research examining the statistical association between survival and mortality in NHANES cohorts (see for example [44, 52, 35, 44]). However, none of these papers examine the uncertainty quantification problems and only focus on examining the statistical association, for example, in terms of measures of the ROC curve such as the area under the curve (AUC) with a pointwise estimation. A basic preliminary exploratory analysis of the NHANES dataset can be found in Figure 5.
Uncertainty quantification can be determinant in making informed public health decisions about the life expectancy of the population. This, in turn, impacts decision-making in multiple settings, such as social policies and telecare in residences—for instance, in rural areas.
Figure 5: Estimated survival functions for four different groups in the NHANES data. Each marginal Turnbull estimator is fed with data from one of the four subsets, created by partitioning the dataset based on whether individuals’ age and TLAC are below or above the medians of these covariates. Note that older individuals who engage in sufficient physical activity exhibit survival probabilities comparable to those of younger individuals who lead sedentary lifestyles.
Our off-the-shelf algorithm in this case is the following Cox model:
$h(T\mid\text{Age},\text{TLAC})=h_{0}(T)\exp(-\text{Age}\beta_{\text{Age}}-\text{TLAC}\beta_{\text{TLAC}})$
where $h(\cdot)$ denotes the baseline hazard function, $T$ is patient survival in months, Age is measured in years and TLAC is a variable related to the physical activity levels of the individual (with higher physical activity levels corresponding to higher TLAC values). We use uncervals with $e=*$. Table 12 shows the intervals $[\text{left},\infty]$ for the 20th and 80th percentiles of the subsample considered for Age and TLAC, indicating the LPB with a probability of 95 percent. Younger individuals (Age=58) and more active (TLAC=3241) can live at least 124 months with high probability. Conversely, more elderly individuals (Age=73) and less active (TLAC=2263) have a threshold of 45.45 months, about one-third of the time. In intermediate cases, there is an increase of 50 percent for increased TLAC values and 100 percent for reducing the Age to 58 years, indicating a stronger influence of Age compared to TLAC. The results indicate that the uncertainty of survival depends on the patients’ characteristics, and informed decisions must take into account public health criteria based on uncertainty. For example, inactive people aged 73 years lose 30 percent survival with respect to the active group.
TLAC $=2263$ TLAC $=3241$
Age $=58$ $(87.89,+\infty)$ $(124.24,+\infty)$
Age $=73$ $(45.45,+\infty)$ $(62.12,+\infty)$
5.2.1 Goodness of fit for conditional survival function for interval-censored data
The goal of this subsection is to illustrate how we can exploit uncervals to evaluate whether using a Cox model in the NHANES case application is well specified via a goodness-of-fit approach. The general idea of extending such a procedure is similar to [17], but in that research the authors focus solely on right-censored data.
Assuming that the random sample $\mathcal{D}_{N}$ is distributed according to a theoretical conditional cdf $F_{0}$, under the null hypothesis $H_{0}:F=F_{0}$ we have $F\left(X_{i},T_{i}\right)\sim U(0,1)$ in virtue of the probability integral transform theorem provided that $F_{0}$ is continuous. Therefore, testing the null hypothesis is equivalent to testing $H_{0}:F\circ F_{0}^{-1}=F_{\text{unif}}$, where $F_{\text{unif}}$ denotes the uniform distribution function on $(0,1)$. Notice that since we have interval-censored data, we do not observe the failure times $T_{i}$ but instead observe $(L_{i},U_{i})$. Under non-informative interval-censoring (Assumption 1), the empirical distribution function $\frac{1}{n}\sum_{i=1}^{n}1\{\Phi^{*}_{i}\leq\cdot\}$ converges in distribution to $U(0,1)$ under the null hypothesis. Recall that
$\Phi^{*}_{i}=\hat{F}_{1}(L_{i},X_{i})+U(0,1)\cdot(\hat{F}_{1}(U_{i},X_{i})-\hat{F}_{1}(L_{i},X_{i})).$
A natural application in our NHANES application case arises, which is testing whether the Cox model is well-specified by evaluating the pseudo-random scores $\Phi^{*}_{i}\sim U(0,1),\ i=1,\dots,n$ via a graphical criterion (as shown in Figure 6). In this case, we can see that the method is well-calibrated and there is no evidence that the conditional survival function class is misspecified with a Cox model. A Kolmogorov-Smirnov test would have to be performed in order to conduct a formal hypothesis test.
Users might notice that using the constructed $\Phi^{*}_{i}$, based on $\hat{F}_{1}$ as Turnbull’s estimator or the recently explained graphical criterion, also seems to be correct and might lead one to believe that it is acceptable to disregard the covariates. This is entirely expected, as Turnbull’s estimator is fully non-parametric and, therefore, will naturally produce residuals with the correct distribution. However, what we are testing is whether the parametric specification formulated in the Cox model is correct for the two chosen covariates: Age and TLAC.
Figure 6: In blue, empirical cumulative distribution function of $\{\Phi^{*}_{i}\}_{i\in\mathcal{I}_{2}}$ using the NHANES database. The ic_sp algorithm (which uses a Cox model) was utilized as the base procedure to compute $\hat{F}_{1}$ For comparison, the black line illustrates the theoretical cumulative distribution function of a uniform distribution $U(0,1)$, which is the identity function.
6 Discussion
The present paper proposes a new framework for uncertainty quantification in interval-censored data. We introduce novel algorithms along with new empirical process tools to examine the theoretical properties of the methods presented. In certain special cases, the methods introduced here are conformal algorithms that exhibit non-asymptotic guarantees of the form $\mathbb{P}(T_{N+1}\in\widehat{\mathcal{C}}_{1-\alpha}(X_{N+1}))\geq 1-\alpha$. These proposed methods are a natural progression of those proposed by [5] for right-censored data, extending them to interval-censored data. In this sense, untervals can be seen as a trade-off between [5] and [55]. Diverse numerical results demonstrate clear advantages in terms of coverage approximation, especially when compared to [5] and the quantile-based approach using raw conditional survival models without calibration.
For the theoretical analysis, we introduce a new class of functions that allows to derive new asymptotic properties of the uncertainty quantification methods presented.
We illustrate the advantages of our proposal over a naive adaptation for interval-censored data of [5] in a clinical problem related to sleep disorders. Additionally, we highlight the application and scientific interest of these methods in quantifying the impact of physical activity on patient survival. We also introduce the core steps to use the interval-censored methodology to do goodness of fit in this setting.
As future work, we suggest extending the proposed framework to handle truncated or double-truncated censored data (see [12]). Additionally, we propose examining the methods introduced here for multivariate interval data. We aim to develop tolerance regions through resampling techniques by exploiting the interval process and its corresponding universal-Donsker property. To the best of our knowledge, this would be the first general framework for tolerance regions [45, 75] for interval-censored data.

### v2

@@ I.1 Simulating an interval-censored response
$F(t,x)=\mathbb{P}\left(T\leq t\mid X=x\right)=\mathbb{P}\left(p\log\left(sT\right)\leq p\log\left(st\right)\mid X=x\right)$
$=\mathbb{P}\left(-r(x)+H\leq p\log\left(st\right)\right)$
$=\mathbb{P}\left(H\leq p\log\left(st\right)+r(x)\right)$
@@
Therefore, Equation 7 also encompasses a Cox model with baseline cumulative hazard $H_{0}(t)=(st)^{p}$, corresponding to a Weibull distribution of scale and shape parameters $s$ and $p$ respectively.
I.2 Verification of asymptotic properties
We prepare four different simulation scenarios, the setup of which is described in Table 1. The underlying true times are simulated according to model 7. These are then censored with a case II interval-censoring mechanism 333see https://cran.r-project.org/web/packages/icenReg/index.html. Time between inspections is distributed as runif(min = 0, max = inspectLength). Then, we check the frequency with which the intervals outputted by our algorithm contain the true time-to-event over 100 test points. We average the results across 100 different training sets. This procedure is repeated for six different nominal coverages ($0.991-\alpha=0.5,0.7,0.8,0.9,0.95,$ 0.99) and four ($N=100,200,500,1000$) different sample sizes. For each row in Table 1, we choose different base estimators for $\hat{F}_{1}$. In the “Linear PH” case, we use ic_sp(Surv(L, U, type = ’interval2’), model = ’ph’) from the icenReg package to fit $\hat{F}_{1}$. In the “Non-linear PH” case, we use Interval Censored Recursive Forests Hunyong Cho & Kosorok (2022) with ntree=20, nfold=2. The marginal case is the same as “Linear PH” but with no covariates specified. For the “Oracle” mode, the true survival function, as derived in LABEL:ors, is directly passed to our approach.
@@ I.2 Verification of asymptotic properties
Figure 8: Averaged over test points vs. nominal coverage for the setup in Section 3.2. The x-axis comprises $\alpha\in\{0.5,0.3,0.2,0.1,0.05,0.01\}$ in all graphs. Sample size refers to $N$. It must be noted that throughout our simulations, uncervals was operating $b=0.5$. This means that we are considering intervals whose image by $\hat{F}_{1}(\cdot,x)$ is centered around 0.5.
I.3 Comparing different LPBs
Table 4: Miscoverage $\cdot 100$ for naive quantile, $\alpha=0.2$
@@ I.3 Comparing different LPBs
Table 5: Miscoverage $\cdot 100$ for uncervals, e=0 , $\alpha=0.2$
@@
Table 6: Miscoverage $\cdot 100$ for uncervals, e=* , $\alpha=0.2$
@@
Table 7: Miscoverage $\cdot 100$ for naive quantile , $\alpha=0.1$
@@
Table 8: Miscoverage $\cdot 100$ for uncervals, e=0 , $\alpha=0.1$
@@
Table 9: Miscoverage $\cdot 100$ for uncervals, e=*, $\alpha=0.1$
@@
Table 10: Miscoverage $\cdot 100$ for naive quantile, $\alpha=0.05$
@@
Table 11: Miscoverage $\cdot 100$ for uncervals, e=0, $\alpha=0.05$
@@
Table 12: Miscoverage $\cdot 100$ for uncervals, e=*,$\alpha=0.05$
@@
I.4 Conditional coverage
Figure 9: Becnhmark plots for for our approach vs naively taking the $\alpha$- th quantile using the empirical square root $L_{2}$ error between the estimated coverage propensity and nominal $(1-\alpha)$ (err). Notice how uncervals performs better most of the times, especially when increasing the sample size to $n=1000$. See plots for $\alpha=0.1$ in the supplement
Figure 10: Boxplots for the results condensated in the tables in the section of the main document with the same title.